OpenStudy (anonymous):

what is the distance between the points (-1,2) and (2,6)?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

apply the distance formula: \[\sqrt{(x2-x1)^2 + (y2-y1)^2}\] where the points given are (x1, y1) and (x2, y2). So in your case, choose one set as x1, y1 (doesn't matter, just keep consistent). so: \[\sqrt{(-1-2)^2 + (2-6)^2} = \sqrt{(-3)^2 + (-4)^2} = \sqrt{25}= 5\]